# Quasi-universal relations in the context of future neutron star detections

### v1

@@ Abstract
The equation of state dependence of neutron star’s astrophysical features modeling is key to our understanding of isospin asymmetric and dense matter. However, there exists a series of almost equation-of-state independent relations reported in the literature, called quasi-universal relations, that are used to determine neutron star radii and moments of inertia from X-ray and gravitational wave signals. Using sets of equations of state constrained by multi-messenger astronomy measurements and nuclear-physics theory, we discuss quasi-universal relations in the context of future gravitational-wave detectors Cosmic Explorer and Einstein Telescope, and X-ray detector STROBE-X. We focus on relations that involve the moment of inertia $I$, the tidal deformability $\Lambda$ and the compactness $C$: $C(\Lambda)$, $I(\Lambda)$ and $I(C)$. The quasi-universal fits and their associated errors are constructed with three different microphysics approaches which include state of the art nuclear physics theory and astrophysical constraints. Gravitational-wave and X-ray signals are simulated with the sensitivity of the next generation of detectors. Equation of state inference on those simulated signals is performed to assess if quasi-universal relations will offer a better precision on the extraction of neutron star’s macroscopic parameters than equation of state dependent relations. We confirm that the relation $I(\Lambda)$ offers a more pronounced universality than relations involving the compactness regardless of the equation of state set. We show that detections with the 3rd generation of gravitational wave detectors and the X-ray detector STROBE-X will be sensitive to the fit error marginalization technique. We also find that the sensitivity of those detectors will be sufficient that using full equation of state distributions will offer better precision on extracted parameters than quasi-universal relations. We also note that nuclear physics theory offers a more pronounced equation of state invariance of quasi-universal relations than current astrophysical constraints.
@@ I Introduction
As the densest stars in the Universe, neutron stars are particularly well suited to investigate ultra-dense matter. Our understanding of the innermost layers of neutron stars remains uncertain due to limitations of nuclear physics laboratories to reach equivalent regimes of temperature and density. Neutron star astrophysical features strongly depend on the equation of state of ultra-dense matter, thus offering the opportunity to probe the neutron star interior with multi-messenger astronomy. On the other hand, a series of relations between various neutron star observables, referred to as quasi (or almost) universal relations, were empirically found to depend weakly on the exact equation of state [1].
@@
While the perfectly holding universality (as per the no-hair theorem) of isolated and stationary black holes in the gravitional theory of General Relativity is in principle not applicable to compact stars, their external gravitational field presents features that are almost independent of the neutron star’s interior. This (almost) universality applies to non-magnetized neutron stars on a static metric in General Relativity and holds for magnetized [2] and spinning neutron stars [3], as well as in modified gravity theories [4]. The underlying physics of this universality has been connected to approximate no-hair relations, where high order multiple moments of compact stars are approximately determined by low order multipole moments, and to the self-similarity of isodensities in compact star. For a detailed introduction and history of universality in compact stars, as well as the derivation of the approximate no-hair relations, details on the I-Love-Q relations and also on the black-hole limit of this almost universality, we refer to the extensive review and work of Ref. [5] and references therein.
@@
The era of multi-messenger astronomy has allowed for the detection of various neutron star’s astrophysical features and is expected to provide increasingly more precise observational data in the future. Almost equation of state independent relations between macroscopic parameters of neutron stars are well suited to extract the astrophysical detection of one parameter from the measure of another (or others). The measurement of post-Keplerian parameters in binary systems via pulsar timing has provided the most precise measurements of neutron star’s gravitational mass $M$. Various wavelength of the electromagnetic spectrum (radio, X-ray and optics) as well as gravitational wave signals from binary neutron star mergers have provided a large number of mass measurements [^1] For a list of neutron star masses measurements with reported precision, see https://compose.obspm.fr/resources.. The Neutron star Interior Composition Explorer (NICER) telescope, which relies on the effects of General Relativity on the detection of hot spots on the surface of a rotating neutron star, has provided the simultaneous measurement of both the gravitational mass and the radius $R$ for two sources (J0030$+$0451 and J0740$+$6620) [6, 7, 8, 9]. Although the precision of the NICER telescope has made some constraint on the equation of state of dense matter, the future Spectroscopic Time-Resolving Observatory for Broadband Energy X-rays (STROBE-X) [10] is expected to offer measurements of mass-radius contours two to three times tighter. The tidal deformability $\Lambda$ of a neutron star was for the first time from the gravitational wave measurement of the double neutron star binary merger source GW170817 [11]. The significant increase in sensitivity brought forth by the next generation of gravitational wave detectors, such as the Cosmic Explorer [12] and the Einstein Telescope [13, 14], is expected to increase the number and precision of detections by several orders of magnitude. Double pulsar binaries are the most promising systems to detect the star’s moment of inertia $I$; the famous double pulsar PSR J0737073707370737-$3039$ for which more than 15 years of data was gathered [15] has not permitted the direct measurement of the moment of inertia yet. However, Ref. [16] have been able to use a quasi-universal relation that involves $I$, $M$ and $\Lambda$ to extract constraints on the moment of inertia of J0737--3039A.
@@
In this paper, we assess the quasi-universality of three relations and discuss their use in the context of next generation of detectors. In Section II, the quasi-universal relations used in this paper and the equation of state sets on which they are based are discussed. We then give details on quasi-universal relations fits and how to introduce the equation of state variability as the fit’s error. We show how next-generation detections are simulated. In Section III, we compare the quasi-universal relations designed with different equation of state sets. We then discuss the impact of different marginalizations of the fit error on the extraction of neutron star’s macroscopic parameters. We contrast the detector sensitivity with the equation of state variability of quasi-universal relations. Finally, we use equation of state inference to compare the accuracy of parameter extraction with quasi-universal relations and equation of state dependent relations. In the Appendix, we present the mass, radius, tidal deformability and moment of inertia modeling, as well as parameters for the quasi-universal relation fits.
@@ II Methods
In this paper, we explore astrophysical features of a neutron star within the theory of General Relativity, assuming the line element of a spherically symmetric, static and isotropic space-time. The mass $M$ and radius $R$ are solved using the Tolman-Oppenheimer-Volkov [17, 18] differential equations. The moment of inertia $I$ is determined in the slow-rotation approximation of Ref. [19] and the tidal deformability follows the Ref. [20] derivation. For details on the modeling of macroscopic parameters, see Appendix A.
II.1 Quasi-universal relations
II.1.1 $C(\Lambda)$, $I(\Lambda)$ and $I(C)$
In this paper, we study the following relations:
@@ II.1.1 $C(\Lambda)$, $I(\Lambda)$ and $I(C)$
    The relation between the compactness and the dimensionless tidal deformability $C(\Lambda)$, with ${C=GM/(Rc^{2})}$. In the assumption that this relation is equation of state independent, it can be used to extract the radius of the neutron star from the gravitational wave measurement of its mass and tidal deformability, see e.g., Ref. [21].
@@
    The relation between the dimensionless moment of inertia and the compactness $\bar{I}(C)$. In the assumption that this relation is equation of state independent, it can be used to extract the moment of inertia from the simultaneous measurement of the mass and radius by telescope such as NICER or Strobe-X, see e.g., Ref. [22].
Not all of the above mentioned relations are equal in the face of equation of state independence (see Ref. [5, 23]), but they are all considered in the literature to be quasi-universal. If the relations hold through equation of state variation, it is also possible to parameterize them with what is referred to as fits, see, e.g., Refs. [24, 1, 5, 25, 26, 27] or Section II.3.2 of the present paper.
II.1.2 Quasi-universal or data driven relations
The universality qualifier attributed to the relations studied in this paper should relate astrophysical features of neutron star’s in spite of the description of the neutron star interior, or in other words the Equation of State (EoS). It should be possible to use minimally constrained EoSs (only causal and themodynamically consistent) and still retain the quasi-universality in the relations. Such relations would reflect the EoS independence driven by General Relativity, not a decreased EoS variability related to our increasing knowledge on the behavior of ultra-dense matter, or our bias in choosing certain EoSs. The introduction of astrophysical and nuclear constraints on the EoS can appear as stronger universality but would be better described as a data driven narrowing of the EoS dependence.
In this paper, we do not use quasi-universal relations established from minimally constrained EoS sets, but data-driven relations from informed sets. We consider both up to date astrophysical data and nuclear physics data.
@@ II.2.1 Nuclear physics based set
The first set comprises sixty one EoSs established from full nuclear physics calculations. The EoSs are gathered from Ref. [28] to which we add seven EoSs presented in Ref. [29]. They describe cold and catalyzed $\beta$-equilibrated matter as such thermodynamic conditions are relevant for isolated ”adult” (not proto) neutron stars and for the inspiral phase of a neutron star merger [^2] We acknowledge that a global thermodynamic equilibrium (catalyzed) neutron star crust may not be appropriate to describe some of the potential STROBE-X sources located in accreting binaries, but nevertheless we reasonably neglect the impact of an accreted crust on the modeling of macroscopic parameters. .
The EoSs of this set were computed with two different nuclear physics approaches: relativistic mean field theory (x39), Skyrme force energy density functionals (x22). Among the relativistic mean field models are included ten EoSs with a hyperonic core (for a recent review on hyperonic compact stars, see Ref. [30]) and fourteen hybrid models with a core quark phase transition (see, e.g., Refs. [31, 32]); the rest have nucleonic cores.
@@ II.2.2 Agnostic sets
The two other sets are based on an agnostic approach, that is to say that the EoSs are not constructed with specific nuclear physics calculations. The point of such sets is to explore the parameter space of pressure and density and to go beyond our usual EoS constructions. In the following, we discuss two agnostic sets: one parametric which we will refer to as the ”meta-model” set, and one non parametric, later on referred to as the ”Gaussian process” set.
@@
The first of the agnostic sets is based on Ref. [37]. To construct one EoS, the nuclear empirical parameters that are the saturation density, the energy of symmetric matter at saturation density, the symmetry energy, the isoscalar and isovector incompressibility, skewness and kurtosis, the nucleon effective mass and the effective mass isosplit are randomly thrown in intervals determined by nuclear physics laboratory experiments. Taking advantage of chiral-effective-field theory calculations presented in Ref. [38] for pure neutron matter, the $\beta$-equilibrated EoSs can be constrained. For a given collection of nuclear empirical parameters, the low density part of the EoS is reconstructed according to the meta-modeling approach discussed in Ref. [39], with a compressible liquid drop model for the inhomogeneous crust. The high density part is constructed with five polytropes for which the adiabatic index and the polytropic constant are randomly thrown. The meta-model set, respects thermodynamic consistency, sound speed causality and provides EoSs in accordance with nuclear physics laboratory experiments while keeping the freedom permitted by the unknown core behavior and the error bars of nuclear experiments at low density. It is a parametric approach, that is to say it follows a specific functional, which implements bias. This set is constrained by the pulsar timing mass measurement of the millisecond pulsar J0740$+$6620, it is composed of $5\times 10^{4}$ EoSs, and is denoted MM+$\chi$+PSR.
@@
The second agnostic set is publicly available and constructed with the Gaussian process approach discussed in Ref. [40]. Contrary to a parametric approach, the Gaussian process EoSs are not bound by the functional chosen to parameterize, and therefore avoid the sort of bias inherent to parametrized constructions, e.g., the discontinuous sound speed of piecewise polytropes in the meta-model set. The Gaussian process set is trained on fifty nuclear physics based EoSs; the selection of EoSs constitutes a bias, but the training is believed to be sufficiently loose that this bias is tamed. The low density part of the EoS is conditioned on the DH crust [41]. This set does not offer a well constrained and well constructed crust and core-crust transition, contrary to the meta-model set and the nuclear physics based set. It follows thermodynamic consistency and sound speed causality. This set of EoSs is constrained by the mass of J0740$+$6620, NICER measurements and by the tidal deformability measurement of GW170817, we refer to it as GP+astro.
@@ II.3.1 Parameter extraction from future astrophysical detections
We refer to parameter extraction as the technique that consists in determining the value of an astrophysical parameter from the measurement of one or several others. To do so, one can either use EoS dependent relations, or quasi EoS independent relations: for example, using a mass measurement we can estimate the radius from the EoS dependent relation $M(R)$, or using a mass and tidal deformability measurement we can estimate the radius from the almost EoS independent relation $C(\Lambda)$.
@@
The reason why the relations presented in Section II.1.1 are preferred in parameter extraction is related to the idea that we can count on the universality to retain a tight error bar on the extracted parameter. But in the context of increasing detector precision, we have to assess whether the EoS variability of the relations discussed in this paper overcomes the detector precision. In other words, we shall assess if the quasi-universality or our current state of the art of astrophysics and nuclear physics constrained relations are sufficient for a precise parameter extraction, or if we will have to keep the relations up to date with upcoming EoS constraints.
If the detector precision imposes the use of increasingly data driven relations in parameter extraction, it is also reasonable to question if Section II.1.1 relations are still more relevant for parameter extraction than EoS dependent relations. It is important to note that Section II.1.1 relations require the detection of at least two parameters to estimate a third, which implies at least two detection errors. EoS dependent relations may involve the detection of only one parameter. Therefore, in the context of increasing detection precision, we have to assess if EoS dependent relations that are sufficiently informed by astrophysical measurements might provide tighter error bars on the extracted parameter than source or data driven relations.
@@
To assess those two points, we simulate neutron-star gravitational and X-ray detections and estimate their associated detection errors:
@@
  • A double neutron star binary merger emitting gravitational waves, for which individual masses and tidal deformabilities are recovered with the public parameter estimation software Bilby [42] with ROQ likelihoods [43] suited for third generation telescope analysis [44]. Sky localization parameters of the binary (right ascension and declination) are fixed, or in other words we assume that an electromagnetic counterpart has been found. The spins are aligned, equal and low ($\chi\leq 0.015$) and we consider a binary with a neutron star of mass ${M_{1}=1.5{\;\rm M}_{\odot}}$ and a neutron star of mass ${M_{2}=1.3{\;\rm M}_{\odot}}$. Tidal deformabilities are calculated from the relation $\Lambda(M)$ determined by the unified equation of state model DD2 [45] as presented in Ref. [33]: ${\Lambda_{1}\simeq 470}$ and ${\Lambda\simeq 1055}$. Simulations are performed with different detector sensitivities: projected O4 sensitivity of LIGO and Virgo facilities (denoted GW-O4), and third generation telescopes Cosmic Explorer and Einstein Telescope sensitivity (denoted GW-3G) [^4] The sensitivity curves for CE and ET (ET-D design [46]) used to simulate the signal in this paper can be found at https://dcc.ligo.org/LIGO-T1500293/public. Updated sensitivity ET-D curves were provided in Ref. [14] recently, but conclusions of our paper remain robust.. The signal to noise ratio of this event is quite high ($\simeq 1800$) although not as high as the source presented in Ref. [44].
[…]

### v2

# Quasiuniversal relations in the context of future neutron star detections

@@ Abstract
The equation of state dependence of a neutron star’s astrophysical features is key to our understanding of isospin asymmetric and dense matter. There exists a series of almost equation of state independent relations reported in the literature, called quasiuniversal relations, that are used to determine neutron star radii and moments of inertia from x-ray and gravitational wave signals. Using sets of equations of state constrained by multimessenger astronomy measurements and nuclear-physics theory, we discuss quasiuniversal relations in the context of future gravitational wave detectors Cosmic Explorer and Einstein Telescope, and the Spectroscopic Time-Resolving Observatory for Broadband Energy X-rays. We focus on relations that involve the moment of inertia $I$, the tidal deformability $\Lambda$, and the compactness $C$: $C(\Lambda)$, $I(\Lambda)$ and $I(C)$. The quasiuniversal fits and their associated errors are constructed with three different microphysics approaches which include state of the art nuclear physics theory and astrophysical constraints. Gravitational-wave and x-ray signals are simulated with the sensitivity of the next generation of detectors. Equation of state inference on those simulated signals is compared to determine if it will offer a better precision on the extraction of a neutron star’s macroscopic parameters than quasiuniversal relations. We confirm that the relation $I(\Lambda)$ offers a more pronounced universality than relations involving the compactness regardless of the equation of state set. We show that detections with the third generation of gravitational wave detectors and future x-ray detectors will be sensitive to the fit error marginalization technique. We also find that the sensitivity of those detectors will be sufficient in that using full equation of state distributions leads to significantly better precision on extracted parameters than quasiuniversal relations. We also note that nuclear physics theory offers a more pronounced equation of state invariance of quasiuniversal relations than current astrophysical constraints.
@@ I Introduction
As the densest stars in the Universe, neutron stars are particularly well suited to investigate ultradense matter. Our understanding of the innermost layers of neutron stars remains uncertain due to limitations of nuclear physics laboratories to reach equivalent regimes of temperature and density. Neutron star astrophysical features strongly depend on the equation of state of ultradense matter, thus offering the opportunity to probe the neutron star interior with multimessenger astronomy. On the other hand, a series of relations between various neutron star observables, referred to as quasi (or almost) universal relations, were empirically found to depend weakly on the exact equation of state [1].
@@
While the perfect universality (as per the no-hair theorem) of isolated and stationary black holes in the gravitational theory of general relativity is, in principle, not applicable to compact stars, their external gravitational field presents features that are almost independent of the neutron star’s interior. This (almost) universality applies to nonmagnetized neutron stars on a static metric in general relativity and holds for magnetized (e.g., [2]) and spinning neutron stars (e.g., [3, 4, 5]), as well as in modified gravity theories (e.g., [6]). The underlying physics of this universality has been connected to approximate no-hair relations, where high order multiple moments of compact stars are approximately determined by low order multipole moments (e.g., [7, 8]) , and to the self-similarity of isodensities in compact stars (e.g., [9]). For a detailed introduction and history of universality in compact stars, as well as the derivation of the approximate no-hair relations, details on the I-Love-Q relations and also on the black-hole limit of this almost universality, we refer to the extensive review and work of Ref. [10] and references therein.
@@
Quasiuniversal relations between macroscopic properties of neutron stars are well suited to extract one parameter from the measurement of others. The era of multimessenger astronomy has allowed for the detection of various astrophysical neutron star features and is expected to provide increasingly more precise observational data in the future. The measurement of post-Keplerian parameters in binary systems via pulsar timing has provided the most precise measurements of neutron star’s gravitational mass $M$, see e.g., Refs [11, 12, 13, 14, 15]. Various wavelengths of the electromagnetic spectrum (radio, x-ray and optics), as well as gravitational wave signals from binary neutron star mergers, have provided a large number of mass measurements [^1] For a list of neutron star masses measurements with reported precision, see https://compose.obspm.fr/resources.. The Neutron star Interior Composition ExplorER (NICER) telescope, which relies on the effects of general relativity on the detection of hot spots on the surface of a rotating neutron star, has provided the simultaneous measurement of both the gravitational mass and the radius $R$ for J0030$+$0451 [16, 17] and of $R$ for J0740$+$6620 [18, 19] with an independent mass measurement based on radio pulsar timing [20]. The future Spectroscopic Time-Resolving Observatory for Broadband Energy x-rays (STROBE-X) [21] is expected to offer measurements of mass-radius contours two to three times tighter. The tidal deformability $\Lambda$ of a neutron star was constrained for the first time from the gravitational wave measurement of the double neutron star binary merger source GW170817 [22]. The significant increase in sensitivity of the next generation of gravitational wave detectors, such as the Cosmic Explorer (CE) [23] and the Einstein Telescope (ET) [24, 25, 26], is expected to increase the number and precision of such observations by orders of magnitude. Double pulsar binaries are the most promising systems to detect the star’s moment of inertia $I$; the famous double pulsar PSR J0737--3039 for which more than 15 years of data was gathered [27], has not yet permitted the direct measurement of the moment of inertia. However, Ref. [28] has been able to use a quasiuniversal relation that involves $I$, $M$ and $\Lambda$ to extract constraints on the moment of inertia of J0737--3039A.
@@
In this paper, we assess the quasiuniversality of three relations and discuss their usefulness in the context of next generation of detectors. We compare the precision achieved on the radius and moment of inertia of a source when using quasiuniversal relations to that derived from full equation of state inference. In Sec. II, the quasiuniversal relations used in this paper and the equation of state sets on which they are based are discussed. We then give details on quasiuniversal relations fits and how to introduce a fit error based on the set of equations of state considered. In Sec. III, we compare the quasiuniversal relations designed with different equation of state sets. We simulate next-generation detections, and demonstrate the use of quasiuniversal relations to extract parameters that are not directly measured. We then discuss the impact of different marginalizations of the fit error on the extraction of neutron star’s macroscopic parameters. Finally, we use equation of state inference from the same simulated detections, and compare to the accuracy of parameter extraction with quasiuniversal relations. In the Appendix, we present the mass, radius, tidal deformability and moment of inertia modeling, as well as parameters for the quasiuniversal relation fits.
@@ II Methods
In this paper, we explore astrophysical features of a neutron star within the theory of general relativity. The mass $M$ and radius $R$ are found using the Tolman-Oppenheimer-Volkov (TOV) [29, 30] differential equations closed by an equation of state (EoS). The moment of inertia $I$ is determined in the slow-rotation approximation of Ref. [31] and the tidal deformability follows the quadrupole perturbation derivation of Ref. [32] derivation. For details on the modeling of macroscopic parameters, see Appendix B.
II.1 Quasiuniversal relations
II.1.1 $C(\Lambda)$, $I(\Lambda)$, and $I(C)$
In this paper, we study the following quasiuniversal relations:
@@ II.1.1 $C(\Lambda)$, $I(\Lambda)$, and $I(C)$
    The relation between the compactness and the dimensionless tidal deformability $C(\Lambda)$, with ${C=GM/(Rc^{2})}$. In the assumption that this relation is equation of state independent, it has been used to extract the radius of a neutron star from the gravitational wave measurement of its mass and tidal deformability, see e.g., Ref. [33].
@@
    The relation between the dimensionless moment of inertia and the compactness $\bar{I}(C)$. In the assumption that this relation is equation of state independent, it can be used to extract the moment of inertia from the simultaneous measurement of the mass and radius by telescopes such as NICER or STROBE-X, see e.g., Ref. [34].
Not all of the above mentioned relations are equally equation of state independence (see Ref. [10, 35]), but they are all described in the literature as quasiuniversal. It is also possible to parametrize them using what is referred to as fits, see, e.g., Refs. [36, 1, 10, 37, 38, 39, 40] or Sec. II.3.2 of the present paper. Universality suggests that the relations studied in this paper should relate astrophysical features of neutron stars for any description of the neutron star interior, i.e. for any EoS. It should then be possible to use widely varying EoSs (only causal and thermodynamically consistent) and still retain the quasiuniversality in the relations. Such relations would not change due a decreased EoS variability related to our increasing knowledge on the behavior of ultradense matter, or our bias in considering certain EoSs. However, in practice, introduction of astrophysical and nuclear constraints on the EoS lead to updated fits of the quasiuniversal relations. This reflects a observation-driven narrowing of the EoS range rather than full universality. Observations can include both astrophysical data and nuclear physics data; see, e.g., Refs. [40, 37].
@@ II.2 Equations of state sets
In this paper, we compare quasiuniversal relations established from several observation-informed EoS sets.
@@ II.2.1 Nuclear physics based set
The first set comprises 61 EoSs established from full nuclear physics calculations. The EoSs are gathered from Ref. [41] to which we add seven EoSs presented in Ref. [42]. They describe cold and catalyzed $\beta$-equilibrated matter as such thermodynamic conditions are relevant for isolated ”adult” (not proto) neutron stars and for the inspiral phase of a neutron star merger [^2] We acknowledge that a global thermodynamic equilibrium (catalyzed) neutron star crust may not be appropriate to describe some of the potential STROBE-X sources located in accreting binaries, but nevertheless we reasonably neglect the impact of an accreted crust on the modeling of macroscopic parameters..
The EoSs of this set were computed with two different nuclear physics approaches: relativistic mean field theory (x39) and Skyrme force energy density functionals (x22). Among the relativistic mean field models are included ten EoSs with a hyperonic core (for a recent review on hyperonic compact stars, see Ref. [43]) and 14 hybrid models with a core quark phase transition (see, e.g., Refs. [44, 45]); the rest have nucleonic cores.
@@ II.2.2 Agnostic sets
The two other sets are based on an agnostic approach, that is to say that the EoSs are not constructed with specific nuclear physics calculations. The point of such sets is to explore the parameter space of pressure and density and to go beyond our usual EoS constructions. In the following, we discuss two agnostic sets: one parametric which we will refer to as the ”metamodel” set, and one nonparametric, later on referred to as the ”Gaussian process” set.
@@
The first of the agnostic sets is based on Ref. [49]. To construct one EoS, the nuclear empirical parameters that are the saturation density, the energy of symmetric matter at saturation density, the symmetry energy, the isoscalar and isovector incompressibility, skewness and kurtosis, the nucleon effective mass and the effective mass isosplit are randomly thrown in intervals determined by nuclear physics laboratory experiments. Taking advantage of chiral-effective-field theory calculations presented in Ref. [50] for pure neutron matter, the $\beta$-equilibrated EoSs can be constrained. For a given collection of nuclear empirical parameters, the low density part of the EoS is reconstructed according to the metamodeling approach discussed in Ref. [51], with a compressible liquid drop model for the inhomogeneous crust. The high density part is constructed with five polytropes for which the adiabatic index and the polytropic constant are randomly thrown. The metamodel set, respects thermodynamic consistency, sound speed causality, and provides EoSs in accordance with nuclear physics laboratory experiments while keeping the freedom permitted by the unknown core behavior and the error bars of nuclear experiments at low density. It is a parametric approach, that is to say it follows a specific functional, which implements bias. This set is constrained by the pulsar timing mass measurement of the millisecond pulsar J0740$+$6620; it is composed of $5\times 10^{4}$ EoSs, and is denoted MM+$\chi$+PSR.
@@
The second agnostic set is publicly available and constructed with the Gaussian process approach discussed in Ref. [52]. Contrary to a parametric approach, the Gaussian process EoSs are not bound by the functional chosen to parametrize and therefore avoids the sort of bias inherent to parametrized constructions [53, 54], e.g., the discontinuous sound speed of piecewise polytropes in the metamodel set. The Gaussian process set at high density is trained on fifty nuclear physics based EoSs; the selection of EoSs constitutes a bias, but the training is believed to be sufficiently loose that this bias is tamed. The low density part of the EoS is conditioned on three crust EoSs. Contrary to the metamodel, this set does not enforce nuclear physics constraints on the crust, a connected core-crust transition, or $\chi$EFT constraints. It follows thermodynamic consistency and sound speed causality. This set of EoSs is constrained by the mass of J0740$+$6620, NICER measurements, and by the tidal deformability measurement of GW170817, we refer to it as GP+astro.
The relation between the pressure $P$ and the baryonic density $n_{B}$ for the two agnostic sets is presented in Fig. 7 in Appendix A.
@@ II.3.1 Parameter extraction from future astrophysical detections
We refer to parameter extraction as determining the value of an astrophysical parameter from the measurement of one or several others. To do so, one can either use:
  • Quasiuniversal relation fits: the measurement of two neutron star parameters lets us extract the measurement of a third directly from the relation.
  • EoS inference: the detection of two neutron star parameters is used to additionally constrain the set of EoSs using Bayesian inference, and the constrained EoSs are used to compute the third parameter based on the measured values.
For example, using a mass and tidal deformability measurement, we can estimate the radius from the quasiuniversal relation $C(\Lambda)$, or we can use hierarchical inference to update the set of possible EoS, and use the resulting set to find the radius from the mass measurement.
@@
The relations presented in Sec. II.1.1 are preferred in parameter extraction because of their simplicity of use and the idea that we can count on the universality to retain a tight error bar on the extracted parameter. In previous usage, quasiuniversal relations were broadly equivalent in parameter extraction when compared to direct EoS inference (for example, in GW170817 Ref. [33]). But in the context of increasing detector precision, we have to assess whether the detector precision overcomes the EoS dependence of the relations discussed in this paper. In other words, we study if quasiuniversal relations conditioned on state of the art astrophysics and nuclear physics will be sufficient for a precise parameter extraction in the context of current and future detectors or if EoS-inference with next-generation astrophysical measurements provides significantly better error bars on the extracted parameters.
@@
To assess those two points, we will simulate neutron-star gravitational wave and x-ray detections and estimate their associated errors:
@@
  • A GW170817-like double neutron star binary merger emitting gravitational waves, for which individual masses and tidal deformabilities are recovered. One of the neutron stars of the binary has a mass ${M_{\rm gw}=1.5{\;\rm M}_{\odot}}$ to which the tidal deformability $\Lambda_{\rm gw}\simeq 470$ is determined by the unified equation of state model DD2 [55] as presented in Ref. [46]. We use the precision reported in Ref. [40] as well as parameter estimation operated with the public software Bilby [56] to simulate the error associated to the mass and tidal deformability, along with a bivariate normal distribution peaked on $M_{\rm gw}$ and $\Lambda_{\rm gw}$. We use two detector sensitivities : projected O4 sensitivity of Laser Interferometer Gravitational Wave Observatory (LIGO) and Virgo facilities (denoted GW-O4) and third generation telescopes CE and ET sensitivity (denoted GW-3G) [^4] The sensitivity curves for CE and ET (ET-D design [57]) used to simulate the signal in this paper can be found at https://dcc.ligo.org/LIGO-T1500293/public. Updated sensitivity ET-D curves were provided in Ref. [26] recently, but conclusions of our paper remain robust..
  • An x-ray source with a neutron star mass ${M_{\rm xray}=1.4{\rm\;M}_{\odot}}$ and a radius established from the equation of state model DD2 $R_{\rm xray}\simeq 13.16$ km. We use the precision reported in Ref. [17] for the detector sensitivity denoted NICER and the projected sensitivity reported in Ref. [21] denoted STROBE-X, to simulate a distribution of masses and radii using a bivariate normal distribution peaked on Mxray and $R_{\rm xray}$.